Deduce the dimension of the following:



1.work



2.momentum



3.power



4.frequency




1
Expert's answer
2021-11-22T10:13:55-0500

Dimension of work:


"[W]=[F][d]=\text M^1\text L^1\text T^{-2}·\text L^1=\text M^1\text L^2\text T^{-2}."

Momentum:


"[p]=[m][v]=\text M^1·\text L^1\text T^{-1}=\text M^1\text L^1\text T^{-1}."

Power:


"[P]=[F][v]=\text M^1\text L^1\text T^{-2}·\text L^1\text T^{-1}=\text M^1\text L^2\text T^{-3}."

Frequency:


"[f]=\frac1{[T]}=\text T^{-1}."